OpenStudy (anonymous):

Find the GCF of the variable expression: 14xy^2, 28x^2y

OpenStudy (turingtest):

14xy^2=(2)(7)(x)(y)(y) 28x^2y=(2)(2)(7)(x)(x)(y) they each share one 2, one 7, an x, and a y so the GCF=(2)(7)(x)(y)=14xy

OpenStudy (anonymous):

\[14xy ^{2}, 28x ^{2}y\]

OpenStudy (turingtest):

yup, that's what I did

OpenStudy (anonymous):

thanks got it :)